OpenStudy (deadshot):

How do I solve x= log12 over log4?

OpenStudy (deadshot):

\[x=\frac{ \log12 }{ \log4 }\]

OpenStudy (horotat):

\[x = \frac{ \log 12 }{ \log 4 } = \frac{ \log 3*4 }{ \log 4 } = \frac{ \log 3 + \log 4 }{ \log 4 } = \frac{ \log 3 }{ \log 2*2 }+\frac{ \log 4 }{ \log 4 }=\frac{ \log 3 }{ \log 2 + \log 2} +1 = \frac{ \log 3 }{ 2*\log 2 }+1\] \[= \frac{ 0.47712125472 }{ 0.60205999132 }+1\approx 1.79\]

OpenStudy (deadshot):

Thanks for the help!

OpenStudy (horotat):

your welcome :)